OpenStudy (lilmike234):

How do you solve the following equation for x? logx+log(x+2)=log(3x)

OpenStudy (mathstudent55):

\(\log a + \log b = \log ab\) Apply the above to the left side.

OpenStudy (lilmike234):

So x would be equal to 0 right ?